OpenStudy (anonymous):

Do i use difference of cubes? Factor and Solve 8x^6-y^3

zepdrix (zepdrix):

Yesss you do :) Recognize that the firs term can be written as a perfect cube.

zepdrix (zepdrix):

\[\large\rm 8=2^3,\qquad\qquad x^6=x^{2\cdot3}=(x^2)^3\]

OpenStudy (anonymous):

so would it be (8x^2)^3-(y)^3

OpenStudy (anonymous):

im pretty sure im wrong :p

zepdrix (zepdrix):

Ooo close,\[\large\rm 8x^6-y^3\quad=\quad 2^3(x^2)^3-y^3\quad=\quad (2x^2)^3-y^3\]

OpenStudy (anonymous):

Oh okay, I get it now! Thank you ;)

zepdrix (zepdrix):

cool :)